# Coded Downlink Massive Random Access and a Finite de Finetti Theorem

### v1

@@ IV Conclusions
This paper connects the finite de Finetti theorem and the problem of communicating sources with certain symmetry to a randomly activated subset of users in downlink massive random access. We develop a relative entropy version of the finite de Finetti theorem and provide a bound on the extra cost of using a codebook generated from i.i.d. mixture distributions to send exchangeable sources to a random subset of $k$ users out of a large pool of $n$ users. The extra cost is independent of $n$, and is at most $O(k)$ for general exchangeable sources, $O(\log k)$ for finite alphabet exchangeable sources, and $O(1)$ for uniform i.i.d. sources.
@@ Appendix A Proof of Theorem 1
Instead of analyzing the encoder output entropy of a particular codebook, we bound the output entropy over an ensemble of codebooks, then show the existence of one codebook satisfying the derived entropy bound.
First, in a massive random access scenario where a random subset of $k$ users are activated among a large pool of $n$ users, when the source distribution is exchangeable, it does not matter which subset of $k$ users are activated. So we only need to consider $p(\mathbf{x})$ with $\mathbf{x}$ taking values on $\mathcal{X}^{k}$.
Let $p(\mathbf{x})$ be an exchangeable source distribution on $\mathcal{X}^{k}$. Let $q(\mathbf{x})$ be an i.i.d. mixture distribution on $\mathcal{X}^{k}$. Let $\mathbf{M}$ be a random codebook generated according to $q(\mathbf{x})$. Define $T=f_{\mathbf{M}}(\mathbf{X},\mathbf{A})$ to be the encoder output from the following process. We generate a source $\mathbf{x}$ according to $p(\mathbf{x})$. We also generate a random codebook $\mathbf{M}$ according to $q(\mathbf{x})$, then find the index of the first codeword in $\mathbf{M}$ that matches the source $\mathbf{x}$ according to (II-B). We first find an upper bound on $H(T)$.
Once an upper bound on $H(T)$ is obtained, we then argue that there must exist at least one good codebook, whose output entropy across all input $\mathbf{x}\sim p(\mathbf{x})$ is upper bounded by the same bound. This is because mixing increases entropy, so
$\overline{H(f_{\mathbf{m}}(\mathbf{X},\mathbf{A}))}\leq H(f_{\mathbf{M}}(\mathbf{X},\mathbf{A}))=H(T),$ (55)
where $H(f_{\mathbf{m}}(\mathbf{X}))$ is the encoder output entropy for a particular codebook $\mathbf{m}$, and the overline denotes the averaging operation over $\mathbf{m}$. Since there must exist at least one codebook $\mathbf{m}^{*}$ such that
$H(f_{\mathbf{m}^{*}}(\mathbf{X},\mathbf{A}))\leq\overline{H(f_{\mathbf{m}}(\mathbf{X},\mathbf{A}))},$ (56)
@@
$R^{*}\leq H(f_{\mathbf{m}^{*}}(\mathbf{X},\mathbf{A}))\leq H(T).$ (57)
@@
Next, we analyze the distribution of $T$. Since the codewords of a random codebook are generated in an i.i.d. fashion according to the distribution $q(\mathbf{x})$, it follows that when conditioned on a source realization $\mathbf{x}$, the probability that the first match occurs at $T=t$ is a geometric distribution with parameter $q(\mathbf{x})$. Then, it follows that the overall distribution of $T$, across all the sources $\mathbf{x}$, must be a mixture of geometric distributions:
@@
We now bound the entropy of this geometric mixture.
@@ A-A Proof of (17)
A direct calculation of the entropy of a mixture of geometric distributions is nontrivial. To circumvent this, we upper bound $H(T)$ using the fact that
@@
A proof of (59) can be found in [15], where a maximum entropy argument is used. The problem is now reduced to bounding $\mathbb{E}[\log(T)]$. To this end, we use the fact that
$\log(t)<\frac{1}{\ln(2)}\left(\sum_{\ell=1}^{t}\frac{1}{\ell}-\frac{1}{2}\right)\leq\frac{1}{\ln(2)}\sum_{\ell=1}^{t-1}\frac{1}{\ell},\text{ for all }t>1$ (60)
to compute the following bound for $\mathbb{E}[\log(T)]$:
$\begin{split}\mathbb{E}[&\log(T)]=\sum_{t=1}^{\infty}\left(\sum_{\mathbf{x}\in\mathcal{X}^{k}}p(\mathbf{x})(1-q(\mathbf{x}))^{t-1}q(\mathbf{x})\right)\log(t)\\ &=\sum_{\mathbf{x}\in\mathcal{X}^{k},p(\mathbf{x})>0}p(\mathbf{x})q(\mathbf{x})\sum_{t=2}^{\infty}(1-q(\mathbf{x}))^{t-1}\log(t)\\ &\leq\sum_{\mathbf{x}\in\mathcal{X}^{k},p(\mathbf{x})>0}p(\mathbf{x})q(\mathbf{x})\frac{1}{\ln(2)}\sum_{t=2}^{\infty}\sum_{\ell=1}^{t-1}(1-q(\mathbf{x}))^{t-1}\frac{1}{\ell}.\end{split}$ (61)
The next step is to compute the inner double summations. Notice that for $10<\alpha<$ 1 we have that
[…]
where the third equality follows from the fact that
$\frac{d}{d\alpha}\sum_{\ell=1}^{\infty}\frac{\alpha^{\ell}}{\ell}=\sum_{\ell=1}^{\infty}\alpha^{\ell}=\frac{1}{1-\alpha},\text{ for }0<\alpha<1.$ (63)
Substituting (62) with $\alpha=1-q(\mathbf{x})$ into (61) yields
$\mathbb{E}[\log(T)]$$\leq$ $\sum_{\mathbf{x}\in\mathcal{X}^{k},p(\mathbf{x})>0}p(\mathbf{x})q(\mathbf{x})\frac{-\log(q(\mathbf{x}))}{q(\mathbf{x})}$ (64)
$=$ $H(\mathbf{X})+D(p\|q).$
@@
$R^{*}\leq H(\mathbf{X})+D(p\|q)+\log(H(\mathbf{X})+D(p\|q)+1)+1.$ (65)
@@ A-B Proof of (18)
The idea here is to choose an $L$ such that the right-hand expression is easy to analyze. To this end, we define $L$ based on the following partition $\mathcal{D}_{1},\mathcal{D}_{2}\ldots\subseteq\mathcal{X}^{k}$ where
@@
The sets $\mathcal{D}_{\ell}$ can be thought of as a binning of all possible source realizations based on their probability of occurrence with respect to the distribution used to generate the codebook. Since $\mathcal{D}_{1},\mathcal{D}_{2}\ldots$ form a partition, each realization of $\mathbf{x}$ is contained in exactly one of these sets. We can define $L(\mathbf{x})$ as the index of the set containing $\mathbf{x}$.
$L(\mathbf{x})=\ell,\quad\text{s.t.}~{}\mathbf{x}\in\mathcal{D}_{\ell}.$ (68)
It follows that for any $\mathbf{x}\in\mathcal{X}^{k}$ with $p(\mathbf{x})>0$, the index of the set that $\mathbf{x}$ belongs to is $\ell=\left\lceil\log\left(\frac{1}{q(\mathbf{x})}\right)\right\rceil$. This means that $L$ takes on at most $\left\lceil\log\left(\frac{1}{q_{\text{min}}}\right)\right\rceil-\left\lceil\log\left(\frac{1}{q_{\text{max}}}\right)\right\rceil$ values, and therefore has an entropy bounded above as
$H(L)$ $\leq\log\left(\left\lceil\log\left(\frac{1}{q_{\text{min}}}\right)\right\rceil-\left\lceil\log\left(\frac{1}{q_{\text{max}}}\right)\right\rceil\right)$
$\leq\log\left(\log\left(\frac{q_{\text{max}}}{q_{\text{min}}}\right)+1\right).$ (69)
@@
$H(T|L=\ell)$ $=-\sum_{t=1}^{\infty}r_{\ell}(t)\log\left({r_{\ell}(t)}\right)$
$\leq-\sum_{t=1}^{\infty}r_{\ell}(t)\log\left({g(t;2^{-\ell})}\right),$ (71)
where $g(t;\theta)$ is a geometric distribution with parameter $\theta$ and the second line follows from the fact that $-\sum_{t}r_{\ell}(t)\log\left(r_{\ell}(t)\right)\leq-\sum_{t}r_{\ell}(t)\log\left(r^{\prime}(t)\right)$ for any distribution $r^{\prime}(t)$. We can now bound $H(T|L)$ using (A-B)
$H(T|L)$ $=\sum_{\ell}^{\infty}\mathrm{Pr}(L=\ell)H(T|L=\ell)$
$\leq-\sum_{\ell=1}^{\infty}\sum_{\mathbf{x}\in\mathcal{D}_{\ell}}p(\mathbf{x})\sum_{t=1}^{\infty}r_{\ell}(t)\log\left({g(t;2^{-\ell})}\right)$
$=-\sum_{\mathbf{x}}p(\mathbf{x})\sum_{t=1}^{\infty}g(t;q(\mathbf{x}))\log\left({g(t;q(\mathbf{x}))}\right)$
$\qquad+\sum_{\ell=1}^{\infty}\sum_{\mathbf{x}\in\mathcal{D}_{\ell}}p(\mathbf{x})D(g(t;q(\mathbf{x}))\|g(t;2^{-\ell})),$ (72)
where the third line follows from direct computation. Note that right-hand expression of (A-B) is the sum of two terms. The first is a weighted sum of the entropy terms of geometric random variables, whereas the second is a weighted sum of relative entropy between geometric distributions. Both terms can be bounded above by exploiting properties of geometric distributions. In particular, for the first term, we use the fact that the entropy of a random variable with distribution $g(t;\theta)$ is upper bounded by $\log(1/\theta)+\log(e)$ bits. Thus, the first term is upper bounded by $\log(1/q(\mathbf{x}))+\log(e)$. For the second term, it can be readily shown that $D(g(t;q(\mathbf{x}))\|g(t;2^{-\ell}))\leq 1$, for $2^{-\ell}\leq q(\mathbf{x})\leq 2^{-\ell+1}$. Hence, it follows that
$\sum_{\ell=1}^{\infty}\sum_{\mathbf{x}\in\mathcal{D}_{\ell}}p(\mathbf{x})D(g(t;q(\mathbf{x}))\|g(t;2^{-\ell}))\leq 1.$ (73)
Putting it all together, we have
$H(T|L)$ $\leq-\sum_{\mathbf{x}}p(\mathbf{x})\left(\log\left({q(\mathbf{x})}\right)+\log(e)\right)+1$
@@
Finally, combining (57), (66), (A-B), and (A-B) yields
$R^{*}\leq H(\mathbf{X})+D(p\|q)+\log\left(\log\left(\frac{q_{\text{max}}}{q_{\text{min}}}\right)+1\right)+3.$ (75)
@@ Appendix B Proof of Theorem 2
From the urn codebook construction, any $k$ distinct entries in a codeword can be represented by a tuple $\mathbf{Y}=(X_{W_{1}},\cdots,X_{W_{k}})$, where $\mathbf{W}=(W_{1},\cdots,W_{k})$ is i.i.d. over $[k]$. Now, if the sampling pattern were collision-free, it would have generated sequences that are distributed according to the original $p({\mathbf{x}})$. Define the set of all collision-free $\mathbf{w}$ as: =$\mathcal{F}=\{\mathbf{w}\in[k]^{k}\ |\ w_{i}\neq w_{j}\text{ for all }i\neq j\}$. We have
$q(\mathbf{x})$ $=\sum_{\mathbf{w}\in[k]^{k}}\mathrm{Pr}(\mathbf{W}=\mathbf{w})\mathrm{Pr}(\mathbf{Y}=\mathbf{x}|\mathbf{W}=\mathbf{w})$
$\geq\sum_{\mathbf{w}\in\mathcal{F}}\mathrm{Pr}(\mathbf{W}=\mathbf{w})\mathrm{Pr}(\mathbf{Y}=\mathbf{x}|\mathbf{W}=\mathbf{w})$
$=\sum_{\mathbf{w}\in\mathcal{F}}\frac{1}{k^{k}}p(\mathbf{x})=\frac{k!}{k^{k}}p(\mathbf{x}).$ (76)
where in the third line we used the fact that $\mathrm{Pr}(\mathbf{W}=\mathbf{w})=\tfrac{1}{k^{k}}$ since all sampling patterns are equally likely. Noting that $k!>k^{k}e^{-k}$, we have $\log\left(\tfrac{k!}{k^{k}}\right)<k\log(e)$. This implies that $D(p\|q)=\sum p(\mathbf{x})\log\tfrac{p(\mathbf{x})}{q(\mathbf{x})}\leq k\log(e)$.
@@
Next, we show that $D(p\|q)\leq|\mathcal{X}|\log(k+1)$. This bound is a consequence of the method of types and its relation to exchangeability. For vector $\mathbf{x}\in\mathcal{X}^{k}$, let $\mathcal{T}_{\mathbf{x}}^{(k)}=\{\mathbf{s}\in\mathcal{X}^{k}\ |\ \hat{p}_{\mathbf{s}}=\hat{p}_{\mathbf{x}}\}$ denote its type class, where the notation $\hat{p}_{\mathbf{s}}$ is taken from (23). Due to the exchangeability of $p(\mathbf{x})$, if $\mathbf{x}^{(1)},\mathbf{x}^{(2)}\in\mathcal{T}_{\mathbf{x}}^{(k)}$ then $p(\mathbf{x}^{(1)})=p(\mathbf{x}^{(2)})$. The idea is to lower bound $q(\mathbf{x})$ by restricting attention to the sequences within $\mathcal{T}_{\mathbf{x}}^{(k)}=\{\mathbf{s}\in\mathcal{X}^{k}\ |\ \hat{p}_{\mathbf{s}}=\hat{p}_{\mathbf{x}}\}$:
$q(\mathbf{x})$ $=\sum_{\mathbf{s}\in\mathcal{X}^{k}}p(\mathbf{s})\left[\prod_{i=1}^{k}\hat{p}_{\mathbf{s}}(x_{i})\right]$
$=\sum_{\mathbf{s}\in\mathcal{X}^{k}}p(\mathbf{s})2^{-k(H(\hat{p}_{\mathbf{x}})+D(\hat{p}_{\mathbf{x}}\|\hat{p}_{\mathbf{s}}))}\geq\sum_{\mathbf{s}\in\mathcal{T}_{\mathbf{x}}^{(k)}}p(\mathbf{x})2^{-kH(\hat{p}_{\mathbf{x}})}$
$=|\mathcal{T}_{\mathbf{x}}^{(k)}|p(\mathbf{x})2^{-kH(\hat{p}_{\mathbf{x}})}\geq\frac{1}{(k+1)^{|\mathcal{X}|}}p(\mathbf{x}),$ (77)
where the second line follows from Theorem 11.1.2 in [16] and the last line follows from the fact that $|\mathcal{T}_{\mathbf{x}}|\geq\frac{1}{(k+1)^{|\mathcal{X}|}}2^{kH(\hat{p}_{\mathbf{x}})}$ [16]. This implies that $D(p\|q)\leq|\mathcal{X}|\log(k+1)$.

### v4

@@ IV Conclusions
This paper connects the finite de Finetti theorem and the problem of communicating sources with certain symmetry to a randomly activated subset of users in downlink massive random access. We develop a KL divergence version of the finite de Finetti theorem and provide a bound on the extra cost of using a codebook generated from an i.i.d. mixture distribution to encode and send exchangeable sources to a random subset of $k$ users out of a large pool of $n$ users. The extra cost is independent of $n$, and is at most $O(k)$ for general exchangeable sources, $O(\log k)$ for finite alphabet exchangeable sources, and $O(1)$ for uniform i.i.d. sources.
V Acknowledgment
We thank the anonymous reviewers for their comments which have improved the presentation of this paper and simplified the proof of Theorem 1.
@@ Appendix A Proof of Theorem 1
First of all, in a massive random access scenario where a random subset of $k$ users become active among a large pool of $n$ users, when the sources are exchangeable, it does not matter which subset of $k$ users are activated. So we only need to consider $p(\mathbf{x})$ with $\mathbf{x}$ taking values on $\mathcal{X}^{k}$.
Instead of constructing a specific codebook, the main idea of the proof is to bound the entropy of the encoder output over an ensemble of random codebooks. Let $p(\mathbf{x})$ be an exchangeable distribution on $\mathcal{X}^{k}$. Let $q(\mathbf{x})$ be an i.i.d. mixture distribution on $\mathcal{X}^{k}$. Define $T=g_{\mathbf{M}}(\mathbf{X},\mathbf{A})$ to be the encoder output from the following process. We generate sources $\mathbf{x}$ according to $p(\mathbf{x})$. We also generate an infinite-size random codebook $\mathbf{M}$ according to $q(\mathbf{x})$. We find $T$, the index of the first codeword in $\mathbf{M}$ that matches the sources $\mathbf{x}$, according to (II-B). We will find an upper bound on $H(T)$.
If the encoder and the decoders share unlimited common randomness, the upper bound on $H(T)$ would have already given us the desired achievability result. In case that common randomness is not available, the encoder and the decoders must share a particular infinite-size codebook. In this case, based on the upper bound on $H(T)$, we can argue that there must exist at least one good codebook, whose output entropy across all input $\mathbf{x}\sim p(\mathbf{x})$ is bounded by the same upper bound. This is because mixing increases entropy, so
$\overline{H(g_{\mathbf{m}}(\mathbf{X},\mathbf{A}))}\leq H(g_{\mathbf{M}}(\mathbf{X},\mathbf{A}))=H(T),$ (80)
where $H(g_{\mathbf{m}}(\mathbf{X}))$ is the encoder output entropy for a particular codebook $\mathbf{m}$, and the overline denotes the averaging operation over $\mathbf{m}$. Since there must exist at least one codebook $\mathbf{m}^{*}$ such that
$H(g_{\mathbf{m}^{*}}(\mathbf{X},\mathbf{A}))\leq\overline{H(g_{\mathbf{m}}(\mathbf{X},\mathbf{A}))},$ (81)
@@
$R^{*}<H(g_{\mathbf{m}^{*}}(\mathbf{X},\mathbf{A}))+1\leq H(T)+1.$ (82)
@@
Next, we analyze the distribution of $T$. Since the codewords of a random codebook are generated in an i.i.d. fashion according to the distribution $q(\mathbf{x})$, it follows that when conditioned on the sources $\mathbf{x}$, the probability that the first match occurs at $T=t$ is a geometric distribution with parameter $q(\mathbf{x})$. Then, it follows that the overall distribution of $T$, across all the sources $\mathbf{x}$, must be a mixture of geometric distributions:
@@
We proceed with bounding the entropy of this geometric mixture.
@@ A-A Proof of (25)
An exact calculation of the entropy of a mixture of geometric distributions is laborious. To circumvent this, we upper bound $H(T)$ using the fact that
@@
A proof of (84) can be found in [18], where a maximum entropy argument is used. The problem is now reduced to bounding $\mathbb{E}[\log(T)]$.
To do so, we apply the conditional version of Jensen’s inequality:
$\mathbb{E}[\log(T)|\mathbf{X}=\mathbf{x}]$$\leq$ $\log\left(\mathbb{E}[T|\mathbf{X}=\mathbf{x}]\right)$ (85)
$=$ $\log\left(\frac{1}{q(\mathbf{x})}\right),$ (86)
since for a fixed $\mathbf{X}=\mathbf{x}$, the distribution of $T$ is a geometric distribution with success probability $q(\mathbf{x})$. Averaging over $\mathbf{X}$, we have
$\mathbb{E}[\log(T)]$ $=$ $\sum_{\mathbf{x}\in\mathcal{X}^{k}}p(\mathbf{x})\mathbb{E}[\log(T)|\mathbf{X}=\mathbf{x}]$ (87)
$\leq$ $\sum_{\mathbf{x}\in\mathcal{X}^{k}}p(\mathbf{x})\log\left(\frac{1}{q(\mathbf{x})}\right)$ (88)
$=$ $H(\mathbf{X})+D(p\|q).$ (89)
@@
$R^{*}<H(\mathbf{X})+D(p\|q)+\log(H(\mathbf{X})+D(p\|q)+1)+2.$ (90)
An alternative proof for the above can be obtained by direct computation; see [2].
@@ A-B Proof of (26)
The idea here is to choose an $L$ such that the right-hand side of the above expression is easy to analyze. To this end, we define $L$ based on the following partition $\mathcal{D}_{1},\mathcal{D}_{2}\ldots\subseteq\mathcal{X}^{k}$ where
@@
Assuming that $q(\mathbf{x})$ is non-degenerate, the sets $\mathcal{D}_{\ell}$ can be thought of as a categorization of all possible sources based on their probability of occurrence with respect to $q(\mathbf{x})$, i.e., the distribution used to generate the codebook. Since $\mathcal{D}_{1},\mathcal{D}_{2}\ldots$ form a partition, each realization of $\mathbf{x}$ is contained in exactly one of these sets. We can define $L(\mathbf{x})$ as the index of the set containing $\mathbf{x}$.
$L(\mathbf{x})=\ell,\quad\text{s.t.}\ \ \mathbf{x}\in\mathcal{D}_{\ell}.$ (93)
It follows that for any $\mathbf{x}\in\mathcal{X}^{k}$ with $p(\mathbf{x})>0$, the index of the set that $\mathbf{x}$ belongs to is $\ell=\left\lceil\log\left(\frac{1}{q(\mathbf{x})}\right)\right\rceil$. This means that $L$ takes on at most $\left\lceil\log\left(\frac{1}{\bar{q}_{\text{min}}}\right)\right\rceil-\left\lceil\log\left(\frac{1}{\bar{q}_{\text{max}}}\right)\right\rceil+1$ values (infinite if $\bar{q}_{\text{min}}=0$), and therefore has an entropy bounded from above as
$H(L)$ $\leq\log\left(\left\lceil\log\left(\frac{1}{\bar{q}_{\text{min}}}\right)\right\rceil-\left\lceil\log\left(\frac{1}{\bar{q}_{\text{max}}}\right)\right\rceil+1\right)$ (94)
$\leq\log\left(\log\left(\frac{\bar{q}_{\text{max}}}{\bar{q}_{\text{min}}}\right)+2\right).$ (95)
@@
$H(T|L=\ell)$ $=-\sum_{t=1}^{\infty}r_{\ell}(t)\log\left({r_{\ell}(t)}\right)$ (97)
$\leq-\sum_{t=1}^{\infty}r_{\ell}(t)\log\left({\mathcal{G}(t;2^{-\ell})}\right),$ (98)
where $\mathcal{G}(t;\theta)$ is a geometric distribution with parameter $\theta$ and the second line follows from the fact that $-\sum_{t}r_{\ell}(t)\log\left(r_{\ell}(t)\right)\leq-\sum_{t}r_{\ell}(t)\log\left(r^{\prime}(t)\right)$ for any distribution $r^{\prime}(t)$. We can now bound $H(T|L)$ using (97)
$H(T|L)$ $=\sum_{\ell=1}^{\infty}\mathrm{Pr}(L=\ell)H(T|L=\ell)$ (99)
$\leq-\sum_{\ell=1}^{\infty}\sum_{\mathbf{x}\in\mathcal{D}_{\ell}}p(\mathbf{x})\sum_{t=1}^{\infty}r_{\ell}(t)\log\left({\mathcal{G}(t;2^{-\ell})}\right)$ (100)
$=-\sum_{\ell=1}^{\infty}\sum_{\mathbf{x}\in\mathcal{D}_{\ell}}p(\mathbf{x})\sum_{t=1}^{\infty}r_{\ell}(t)\log\left(r_{\ell}(t)\right)$
$\qquad+\sum_{\ell=1}^{\infty}\sum_{\mathbf{x}\in\mathcal{D}_{\ell}}p(\mathbf{x})\sum_{t=1}^{\infty}r_{\ell}(t)\log\left(\frac{r_{\ell}(t)}{\mathcal{G}(t;2^{-\ell})}\right)$ (101)
$=-\sum_{\mathbf{x}}p(\mathbf{x})\sum_{t=1}^{\infty}\mathcal{G}(t;q(\mathbf{x}))\log\left({\mathcal{G}(t;q(\mathbf{x}))}\right)$
$\qquad+\sum_{\ell=1}^{\infty}\sum_{\mathbf{x}\in\mathcal{D}_{\ell}}p(\mathbf{x})D(\mathcal{G}(t;q(\mathbf{x}))\|\mathcal{G}(t;2^{-\ell})).$ (102)
Note that right-hand side of the expression of (102) is the sum of two terms. The first is a weighted sum of the entropy terms of geometric random variables, whereas the second is a weighted sum of KL divergences between geometric distributions. Both terms can be upper bounded by using properties of geometric distributions. Using the fact that the entropy of a random variable with distribution $\mathcal{G}(t;\theta)$ is upper bounded by $\log(1/\theta)+\log(e)$ bits, the first term is upper bounded by $\log(1/q(\mathbf{x}))+\log(e)$. For the second term, it can be shown that $D(\mathcal{G}(t;q(\mathbf{x}))\|\mathcal{G}(t;2^{-\ell}))\leq 1$, for $2^{-\ell}\leq q(\mathbf{x})\leq 2^{-\ell+1}$. Hence, it follows that
$\sum_{\ell=1}^{\infty}\sum_{\mathbf{x}\in\mathcal{D}_{\ell}}p(\mathbf{x})D(\mathcal{G}(t;q(\mathbf{x}))\|\mathcal{G}(t;2^{-\ell}))\leq 1.$ (103)
Putting the bounds together, we have
$H(T|L)$ $\leq\sum_{\mathbf{x}\in\mathcal{X}^{k}}p(\mathbf{x})\left(\log\left(\frac{1}{q(\mathbf{x})}\right)+\log(e)\right)+1$
@@
Finally, combining (82), (91), (94), and (A-B) yields
$R^{*}<H(\mathbf{X})+D(p\|q)+\log\left(\log\left(\frac{\bar{q}_{\text{max}}}{\bar{q}_{\text{min}}}\right)+2\right)+4.$ (105)
@@ Appendix B Proof of Theorem 2
As in the proof of Theorem 1, we only need to consider sources $\mathbf{x}\in\mathcal{X}^{k}$ corresponding to any random subset of $k$ active users among the $n$ users.
In the urn codebook construction, we first generate a vector $\mathbf{s}\in\mathcal{X}^{k}$ according to $p(\mathbf{\mathbf{x}})$, then sample the entries of this vector to generate codewords $\mathbf{c}\in\mathcal{X}^{n}$. Note that here we use slightly different notation as in the text and call this vector $\mathbf{s}$.
Our goal is to characterize $q(\mathbf{x})$, the probability that $\mathbf{x}$ appears in any set of $k$ distinct entries in any codeword in the urn codebook. We represent the sampling process by defining the positions of $\mathbf{s}$ that are sampled as $\mathbf{w}=(w_{1},\cdots,w_{k})$ where each $w_{i}$ is i.i.d. over $[k]$. In other words,
$[x_{1},\cdots,x_{k}]^{T}=[s_{w_{1}},\cdots,s_{w_{k}}]^{T}.$ (106)
In the rest of the proof, we use the shorthand “$\mathbf{x}=\mathbf{s}_{\mathbf{w}}$” to denote the above.
@@
Now, if the sampling pattern were collision-free, it would have generated sequences that are distributed according to $p({\mathbf{x}})$ exactly. Define the set of all collision-free $\mathbf{w}$ as
=$\mathcal{F}=\{\mathbf{w}\in[k]^{k}\ |\ w_{i}\neq w_{j}\ \forall i\neq j\}.$ (107)
The actual sampling process may not be collision-free, but we can bound the distribution $q(\mathbf{x})$ as follows:
$q(\mathbf{x})$ $=\sum_{\mathbf{s}\in\mathcal{X}^{k}}\sum_{\mathbf{w}\in[k]^{k}}p(\mathbf{s})p(\mathbf{w})\mathbbm{1}({\mathbf{x}=\mathbf{s}_{\mathbf{w}}})$ (108)
$\geq\sum_{\mathbf{s}\in\mathcal{X}^{k}}\sum_{\mathbf{w}\in\mathcal{F}}p(\mathbf{s})p(\mathbf{w})\mathbbm{1}({\mathbf{x}=\mathbf{s}_{\mathbf{w}}})$ (109)
$=\sum_{\mathbf{w}\in\mathcal{F}}\frac{1}{k^{k}}p(\mathbf{x})$ (110)
$=\frac{k!}{k^{k}}p(\mathbf{x}).$ (111)
where
  • (108) is due to the urn codebook construction process;
  • (109) is due to that collision-free sampling is a subset of all samplings;
  • (110) is due to the facts that under collision-free sampling $p(\mathbf{x})=p(\mathbf{s})$ and that $p(\mathbf{w})=\frac{1}{k^{k}}$ since all sampling patterns are equally likely;
  • (111) is due to the fact that the number of collision-free samplings is $|\mathcal{F}|=k!$.
Noting that $k!>k^{k}e^{-k}$, we have $\log\left(\frac{k^{k}}{k!}\right)<k\log(e)$. This implies that
$D(p\|q)$$\leq$ $\sum_{\mathbf{x}\in\mathcal{X}}p(\mathbf{x})\log\left(\frac{p(\mathbf{x})}{q(\mathbf{x})}\right)$ (112)
$\leq$ $\log\left(\frac{k^{k}}{k!}\right)$ (113)
$<$ $k\log(e).$ (114)
@@
Next, we show that $D(p\|q)\leq|\mathcal{X}|\log(k+1)$. This bound is a consequence of the method of types and its relation to exchangeability. For vector $\mathbf{x}\in\mathcal{X}^{k}$, let its type class be denoted as
$\mathcal{T}_{\mathbf{x}}^{(k)}=\{\mathbf{s}\in\mathcal{X}^{k}\ |\ \hat{p}_{\mathbf{s}}=\hat{p}_{\mathbf{x}}\},$ (115)
where the notation $\hat{p}_{\mathbf{s}}$ is taken from (33).
The idea is to lower bound $q(\mathbf{x})$ by restricting attention to the sequences within $\mathcal{T}_{\mathbf{x}}^{(k)}$:
$q(\mathbf{x})$ $=\sum_{\mathbf{s}\in\mathcal{X}^{k}}p(\mathbf{s})\left(\prod_{i=1}^{k}\hat{p}_{\mathbf{s}}(x_{i})\right)$ (116)
$\geq\sum_{\mathbf{s}\in\mathcal{T}_{\mathbf{x}}^{(k)}}p(\mathbf{s})\left(\prod_{i=1}^{k}\hat{p}_{\mathbf{s}}(x_{i})\right)$ (117)
$=\sum_{\mathbf{s}\in\mathcal{T}_{\mathbf{x}}^{(k)}}p(\mathbf{x})\left(\prod_{i=1}^{k}\hat{p}_{\mathbf{x}}(x_{i})\right)$ (118)
$=|\mathcal{T}_{\mathbf{x}}|\;p(\mathbf{x})\prod_{i=1}^{k}\hat{p}_{\mathbf{x}}(x_{i}),$ (119)
where
  • (116) is due to the urn codebook construction, with each term in the summation being the probability that a codeword $\mathbf{x}$ is generated from some $\mathbf{s}\in\mathcal{X}^{k}$ using its empirical distribution;
  • (118) follows because $p(\cdot)$ is exchangeable, so due to the definition of the type class, $p(\mathbf{x})=p(\mathbf{s}),\forall\mathbf{s}\in\mathcal{T}_{\mathbf{x}}^{(k)}$.
The product distribution in (119) can be shown to be equal to $2^{-kH(\hat{p}_{\mathbf{x}})}$ as follows
$\prod_{i=1}^{k}\hat{p}_{\mathbf{x}}(x_{i})$ $=\prod_{c\in\mathcal{X}}\left(\hat{p}_{\mathbf{x}}(c)\right)^{N_{\mathbf{x}}(c)}$ (120)
$=\prod_{c\in\mathcal{X}}\left(\hat{p}_{\mathbf{x}}(c)\right)^{k\hat{p}_{\mathbf{x}}(c)}$ (121)
$=2^{k\sum_{c\in\mathcal{X}}\hat{p}_{\mathbf{x}}(c)\log\hat{p}_{\mathbf{x}}(c)}$ (122)
$=2^{-kH(\hat{p}_{\mathbf{x}})},$ (123)
where $N_{\mathbf{x}}(c)$ is the number of occurrences of $c$ in $\mathbf{x}$.
Furthermore, by [22, Theorem 11.1.3], the total number of sequences in $\mathcal{T}_{\mathbf{x}}$ must satisfy
$|\mathcal{T}_{\mathbf{x}}|\geq\frac{1}{(k+1)^{|\mathcal{X}|}}2^{kH(\hat{p}_{\mathbf{x}})}.$ (124)
Combining (119), (123) and (124), it follows that
$q(\mathbf{x})\geq\frac{1}{(k+1)^{|\mathcal{X}|}}p(\mathbf{x}).$ (125)
This implies that
$D(p\|q)\leq|\mathcal{X}|\log(k+1).$ (126)
Together, (113), (114) and (126) give (34).